Solve: The Officejet printer can copy Lisa's dissertation in 22 min. The LaserJet printer can copy the same document in 16min. If the two machines work together, how long would they take to copy the dissertation?
Answer by ptaylor(2198)

Let x=amount of time it takes if both machines are working together
Then together the machines copy at the rate of 1/x of the dissertation per min
Officejet copies at the rate of 1/22 of the dissertation per min
Laserjet copies at the rate of 1/16 dissertation per min
Together they copy at the rate of 1/22 + 1/16 of the dissertation per min
So, our equation to solve is:
1/22 + 1/16=1/x multiply each term by 176x
8x+11x=176
19x=176 or
x=176/19 or 9.263 minutes------------------time needed if both machines work together
CK
In (176/19) min, the officejet copies (1/22)(176/19) =8/19 of the dissertation
In (176/19) min, the laserjet copies (1/16)(176/19) =11/19 of the dissertaion
8/19 + 11/19=19/19 or 1 dissertation
Hope this helps---ptaylor